Ten years ago a man was as old as his wife is now. The man is now twice as old as his wife back then. How old are they both now?
Found 2 solutions by mananth, shivang jindal:
Answer by mananth(16946)

let wife be x now
10 years ago she was x-10.
husband was x 10 years ago
so now he is x+10 years
x+10=2(x-10)
x+10=2x-20
2x-x=30
x= 30

Answer by shivang jindal(3)

let wife be x now
10 years ago she was x-10.
husband was x 10 years ago
so now he is x+10 years
x+10=2(x-10)
x+10=2x-20
2x-x=30
x= 30